What is the pH of a solution containing 0.138 mol L-1 NH3 and 0.253 mol L-1 NH4+? Round your answer to 2 decimal places.

Remember you can find KA and/or KB values in your textbook in chapter 15.
Answer #1

• 8 8_ol | NH 4 0:23 |, NH, Weak base Conjugate qued roof -NH3 a such solution is a buffer solution Rond por is geen by pohtaso, por = 4.74 , log 0.93 _ppH = 4.74 +0.26 = 5.00 Si= o pond pH tpon =:14 pH= 14-5 * pH = 9:00